OpenStudy (anonymous):

What is the solution to the following system? -3x - y = -3 9x + 3y = 9 (3, 3) (3, -3) no solution infinite solutions on the line

hero (hero):

y = 3-3x y = (9-9x)/3 y = y 3(3-3x) = 9-9x 9-9x = 9-9x Infinite Solutions